(1 point) Suppose that X is normally distributed with mean 120 and standard deviation 14.

A. What is the probability that X is greater than 140.58? Probability =

B. What value of X does only the top 14% exceed? X =
Answer #1

Solution:

Given that,

\mu = 120

\sigma = 14

A) p ( x > 140.58 )

= 1 - P ( x < 140.58 )

= 1 - P ( ( x - \mu /\sigma ) < ( 140.58 - 120 / 14 ) )

= 1 - P ( z < 20.58/14 )

= 1 - p ( z < 1.47 )

Using z table

= 1 - 0.9292

= 0.0708

Probability = 0.0708   

B ) P(Z > z) = 14%

1 - P(Z < z) = 0.14

P(Z < z) = 1 - 0.14 = 0.86

P(Z < 1.08 ) = 0.86

z = 1.08

Using z-score formula,

x = z * \sigma + \mu

x = 1.08 * 14 + 120 = 135.12

X = 135.12